Grade 11Physical Chemistry

When sulphur ( in the form of S8 ) is heated at temperature T, at equilibrium, the pressure of S8 falls by 30 percent from 1 atm, because S8 (g) is partially converted into S2(g) . Find the value of Kp for this reaction.

Swastik Tripathy
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Please find below the solution to the asked query:

Chemical reaction:   S8 (g) 4 S2 (g)

initial pressure           1 atm          0
Eqbm. Pressure         (1 - a) atm    a atm


At equilibrium a = 29 % = 29/100 = 0.29

Therefore, equilibrium pressure of S8 = (1 - 0.29) = 0.71 atm

and Equilibrium pressure of S2 = 4 x 0.29 = 1.16 atm

So, equilibrium constant : 

Kp = P(S2)^4 / P(S8) = 1.16^4 / 0.71 = 2.55atm